# Learn and Don’t Forget: Adding a New Language to ASR Foundation Models

### v1

@@ 1 Introduction
With the development and release of pre-trained multilingual end-to-end foundation ASR models [11, 12, 13], utilising a multilingual model is another option for low-resource languages. Zhang et al. [14] proposed a smaller universal monolingual output layer shared across languages for high-quality and high-efficiency multilingual ASR. Chen et al. [15] used hierarchical CTC to leverage language identity throughout the entire encoder-decoder network, aiming to improve ASR performance by correctly identifying languages. Qin et al. [16] utilised multilingual and multilevel unit modeling to improve ASR performance on low-resource Tibetan languages.
The challenge in multilingual models lies in expanding them to new languages without compromising the performance of existing ones, particularly when the training data for new languages is limited or unavailable. In NLP, approaches like language-specific adapters and output heads, as seen in [17, 18], enable parameter-efficient fine-tuning but necessitate prior knowledge of the input language. Alternatively, continual lifelong learning, as proposed by [19], combines data from both existing and new languages for ongoing training. Demonstrating its effectiveness in ASR, [20, 21] have expanded language coverage from 15 to 32 and to 66 languages, respectively.

### v2

@@ 1 Introduction
With the development and release of pre-trained multilingual end-to-end foundation ASR models [11, 12, 13], utilising a multilingual model is another option for low-resource languages. Zhang et al. proposed a smaller universal monolingual output layer shared across languages for high-quality and high-efficiency multilingual ASR [14]. Chen et al. used hierarchical CTC to leverage language identity throughout the entire encoder-decoder network, aiming to improve ASR performance by correctly identifying languages [15] . Qin et al. utilised multilingual and multilevel unit modeling to improve ASR performance on low-resource Tibetan languages [16].
The challenge in multilingual models lies in expanding them to new languages without compromising the performance of existing ones, particularly when the training data for new languages is limited or unavailable. In NLP, approaches like language-specific adapters and output heads, as seen in [17, 18], enable parameter-efficient fine-tuning but necessitate prior knowledge of the input language. Alternatively, continual lifelong learning, as proposed by [19], combines data from both existing and new languages for ongoing training. Demonstrating its effectiveness in ASR, [20, 21] have expanded language coverage from 15 to 32 and to 66 languages, respectively. Pham et al. proposed a weight factorization technique to factorize each weight matrix in the network into language dependent and independent factors [22]. When combined with Elastic Weight Consolidation (EWC), this approach allows a multilingual ASR model to expand its learning from an initial 10 languages to 26 languages without catastrophic forgetting [23, 24].